OpenStudy (anonymous):

What is the value of n in the equation 3n + 2(n + 2) = 9n + 12? −4 −2 2 4

OpenStudy (freckles):

do any distributing that is needed

OpenStudy (freckles):

a(b+c)=ab+ac

OpenStudy (freckles):

so 2(n+2)=?

OpenStudy (anonymous):

the answer is 2 nvm i worked it out haha

OpenStudy (anonymous):

thanks anyways

OpenStudy (freckles):

k:)